Four grams of musk oil are required for each bottle of Mink Caress, a very popular perfume made by a small company in western Siberia. The cost of the musk o is $2.10 per gram. Budgeted production of Mink Caress is given below by quarters for Year 2 and for the first quarter of Year 3 Year 2 Year 3 First Second Third Fourth First 108,000 Budgeted production, in bottles 98,000 128,000 188,000 138,000 Musk oil has become so popular as a perfume ingredient that it has become necessary to carry large inventories as a precaution against stock-outs. For this reason, the inventory of musk oil at the end of a quarter must be equal to 20% of the following quarters production needs. Some 78,400 grams of musk oil will be on hand to start the first quarter of Year 2. Required: Prepare a direct materials budget for musk o by quarter and in total, for Year 2. (Round Unit cost of raw materials answers to 2 decimal places.) Mink Caress Direct Materials Budget Year 2 Quarter First Second Third Fourth Year Required production in units of finished goods Units of raw materials needed per unit of finished goods Units of raw materials needed to meet production 0 0 Total units of raw materials needed 0 0 0 0 0 Units of raw materials to be purchased Unit cost of raw materials Cost of raw materials to purchased
Answer #1
A1 B C D E F G H I J K
2 Inventory of Gram Oil in the beginning of Year2 first quarter is 78,400.
3 Ending Inventory required is 20% of next quarter requirement
4 Total materials needed should be calculated by adding the production requirement for the quarter with the ending inventory required for next quarter production.
5 The amount needs to be purchased is calculated by substracting beginning inventory from quantity of materials needed.
6 Year 2 Year3
7 Quarter Quarter
8 First Second Third Fourth First
9 Required Production in units of finished goods 98000 128000 188000 138000 108000
10 Units of raw materials needed per unit of finished goods 4 4 4 4 4
11 Units of raw materials needed to meet production 392000 512000 752000 552000 432000
12 Required Ending Inventory 102400 150400 110400 86400
13 Total units of materials needed 494400 662400 862400 638400
14 Beginning Inventory 78400 102400 150400 110400
15 Units of raw materials to be purchased 416000 560000 712000 528000
16 Unit cost of raw materials $2.10 $2.10 $2.10 $2.10
17 Cost of raw materials to purchased $873,600.00 $1,176,000.00 $1,495,200.00 $1,108,800.00
18

Formula sheet

A1 B C D E F G H I J K
2 Inventory of Gram Oil in the beginning of Year2 first quarter is 78,400.
3 Ending Inventory required is 0.2 of next quarter requirement
4 Total materials needed should be caculated by adding the production requirement for the quarter with the ending invetory required for next quarter production.
5 The amount need to be purchased is calculated by substracting beginning inventory from quanitity of materials needed.
6 Year 2 Year3
7 Quarter Quarter
8 First Second Third Fourth First
9 Required Production in units of finished goods 98000 128000 188000 138000 108000
10 Units of raw materials needed per unit of finished goods 4 4 4 4 4
11 Units of raw materials needed to meet production =E9*E10 =F9*F10 =G9*G10 =H9*H10 =I9*I10
12 Required Ending Inventory =F11*$D$3 =G11*$D$3 =H11*$D$3 =I11*$D$3
13 Total units of materials needed =E11+E12 =F11+F12 =G11+G12 =H11+H12
14 Beginning Inventory =78400 =E12 =F12 =G12
15 Units of raw materials to be purchased =E13-E14 =F13-F14 =G13-G14 =H13-H14
16 Unit cost of raw materials 2.1 2.1 2.1 2.1
17 Cost of raw materials to purchased =E15*E16 =F15*F16 =G15*G16 =H15*H16
18